Determine the value of m so that (x-3) is a factor of 4x^2 – 6x + m.
Answer by vksarvepalli(154)

just use remainder theorem
according to it if f(x) is divisible by x-a then f(a)=0
so here f(x)= 4x^2 – 6x + m
and f(3)=0
so 4*9 - 6*3 +m =0
=> 18+m=0
=> m=-18
ans.